If cos(A-B) = cos 15° and cos(A+B)= cos 75° , the values of A and B. Hence without using a calculator and by using the sum and difference formula, show that cos 15° - cos 75° = 2 sin 45° sin 30°


1
Expert's answer
2022-01-11T00:21:25-0500

"A = 45^o \space \space B=30^o"

"-\begin{cases}
 cos(A-B) = cosAcosB+sinAsinB=cos15^o \\
 cos(A+B) = cosAcosB-sinAsinB=cos75^o
\end{cases}"

"cos15^o-cos75^o=2sinAsinB=2sin45^osin30^o"